• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name E.Chudakov

    Log entry time 13:25:11 on November 26, 2008

    Entry number 250907

    keyword=Meeting today at 16:00. Current status.

    Let us meet today on the 2-nd floor of the counting house at 16:00.

    The current status:
    a) Compton tune yesterday was a success.
    b) The compressor for the dehumidifier has been fixed yesterday.
    c) The target is replaced.
    d) ESR crash today at about 8:00 tripped all the HRSL magnets,
    and required about 4h to recover.
    e) Nearly all reference cell data has been taken, moving to
    the polarized 3He soon.


    A copy of this log entry has been emailed to: transversity@jlab.org